OpenStudy (softballgirl372015):

How you graph y=0.5sinx

sam (.sam.):

y=Asin(x) A= Amplitude

OpenStudy (softballgirl372015):

Okay, I understnad that, but I am confused on how to draw the graph, specifically with the periods.

sam (.sam.):

1 full period will be \(2\pi\)

OpenStudy (whpalmer4):

The graph looks just like \(y=\sin x\) in terms of how frequently it goes up and down. However, instead of going between 1 and -1, it goes between 1/2 and -1/2...

sam (.sam.):